OpenStudy (anonymous):

Tori's 20-ounce glass of soymilk has 15 grams of protein. Brady's 16-ounce glass has 12 grams of protein. What is the ratio of Tori's drink size to Brady's drink size?

hero (hero):

Hint:\[\frac{\text{Tori's drink size}}{\text{Brady's drink size}} = \frac{20}{16}\]

OpenStudy (anonymous):

4:3?

hero (hero):

What? How did you get that? Show your work please....

OpenStudy (anonymous):

Ehhh. Idk, I just guessed.

hero (hero):

You can't just guess. You have to divide the numerator and the denominator by the same factor. What factor is common to both 20 and 16?

OpenStudy (anonymous):

4

hero (hero):

Okay, so divide \[\frac{20 \div 4}{16 \div 4}\]

OpenStudy (anonymous):

5/4

OpenStudy (anonymous):

THANKS:D